# Interpretable Responsibility Sharing as a Heuristic for Task and Motion Planning

### v1

@@ 3.2 Multi-Bound Tree Search
The discrete elements of the LGP formulation, detailed in Equation 1, give rise to a decision tree populated with sequences of symbolic states originating from $s_{0}$. The leaf nodes where $s\in S_{\text{goal}}(g)$ are identified as potential solutions, alternatively goal states $s_{g}$. The feasibility of each pathway is assessed by solving the NLP produced by the state sequence from the root to the evaluated node. This process, however, can be computationally demanding as the number of NLPs to consider is typically large. To address this challenge, Multi-Bound Tree Search (MBTS) initially resolves more relaxed versions of Equation 1  [65]. The feasibility of these relaxed instances is a prerequisite for the feasibility of the corresponding original NLP, thus acting as computational lower bounds. For this work, we utilize two bounds: $P_{seq}$ (joint optimization of the mode-switches of the sequence) and $P_{path}$ (the full motion planning problem).
@@ 3.3 Counterfactual Scenarios & Individual Treatment Effect (ITE)
A counterfactual scenario is the scenario that would have occurred if the sample had received a different treatment, which refers to the action applied to a sample  [66, 67]. In binary treatment scenarios, $Y^{\prime}$ represents the counterfactual outcome, simplified as $Y^{\prime}=Y(W=1-w)$, where $w$ is the actual treatment received by the sample. Starting from the same initial conditions and using a fixed planner, only one plan can be generated if transition dynamics are deterministic. This is because deterministic dynamics ensure a single, predictable outcome for a given set of initial conditions and planner. In this study, to understand the effects of incorporating auxiliary objects, we generate counterfactual plans, $P^{\prime}$. These counterfactual plans allow us to observe how the inclusion of different auxiliary objects might alter the outcomes, providing insights into the impact and utility of these auxiliary elements in the planning process.
The treatment effect quantifies the impact of a treatment or intervention on an outcome of interest. In our work, we investigate whether incorporating auxiliary objects will increase the performance of the agent  [66, 67]. As each scenario is unique, we employ Individual Treatment Effect (ITE) to measure the difference in outcomes for a specific sample if it were to receive the treatment versus not receiving it (i.e., shared responsibility or not).
@@
In a non-deterministic setting, $Y_{i}$ refers to the expected value of an outcome considering a single sample. In a deterministic setting, $Y_{i}$ can represent any measurement function of the outcome, again considering a single sample.
@@ 3.4 Rule-Based Representation Learner (RRL)
RRL  [25] is designed to merge scalability with interpretability by employing non-fuzzy rules for data representation and classification. This model introduces a novel method called Gradient Grafting to train discrete models effectively. RRL’s hierarchical model structure consists of layers for feature discretization, rule-based representation, and rule importance evaluation. The improved logical activation functions in RRL convert multiplications into additions using logarithms, thereby avoiding the vanishing gradient problem often encountered in neural networks.
The binarization layer in RRL divides continuous features into bins, allowing for end-to-end feature discretization. Gradient Grafting, the key training method, utilizes gradient information at both discrete and continuous parametric points, optimizing the training process for discrete models. The Gradient Grafting equation as provided in  [25] captures how the gradient updates are adjusted during the training process to accommodate both continuous and discrete parameter states effectively:
@@
where $I(\cdot)$ is an indicator function. This equation determines the effective gradient, $\hat{g}$, used for parameter updates during the training process. It incorporates a logical indicator function that selects between gradients computed for continuous parameters, $g_{\text{continuous}}$, and discrete parameters, $g_{\text{discrete}}$, based on the output of a logical activation function $\sigma_{\text{logical}}$. If the activation function outputs a value greater than or equal to 0.5, the gradient for continuous parameters is used; otherwise, the gradient for discrete parameters is used.
This method facilitates the seamless integration of learning across discrete and continuous parameter spaces, enhancing the training process’s efficiency and effectiveness. It allows the RRL model to leverage the interpretative benefits of rule-based learning while maintaining the adaptability and robustness provided by gradient-based optimization techniques.
@@ 3.5 Correlation and Order-Aware Rule Learning (CARL)
CARL  [26] captures deeper semantic information in rules by ensuring relations are aware of each other and paying attention to logical sequence sensitivity. The model leverages the semantic consistency between the rule body and rule head as its learning objective. The Correlation Module within CARL explores internal correlations between domain relations using a multi-head attention mechanism to select “active relations”, which reduces computational complexity.
Path sampling in CARL involves a random walk-based sampler that gathers closed paths from the knowledge graph, forming probable rules. Positional encoding generates unique positional vectors for each relation using trigonometric functions, distinguishing rules based on their positions within the rule body. The Semantic Learner component iteratively merges relations to reduce the length of the rule body while continuously introducing more semantic information.
The positional encoding equations in CARL are designed to assign a unique encoding to each relation in a rule based on its position, which is crucial for maintaining the order sensitivity in logical sequences.
$PE_{\text{pos},2i}=\sin\left(\frac{\text{pos}}{10000^{\frac{2i}{d_{\text{model}}}}}\right)$ (4)
$PE_{\text{pos},2i+1}=\cos\left(\frac{\text{pos}}{10000^{\frac{2i}{d_{\text{model}}}}}\right)$ (5)
These functions, sine and cosine, are utilized to generate positional vectors, where pos represents the position of a relation in the rule, and $d_{\text{model}}$ is the dimensionality of the model. These vectors aid in distinguishing the roles of the same relation appearing at different positions within the rule sequence, ensuring that the model captures the order-sensitive information effectively.
The Multi-Head Attention mechanism in the Semantic Learner component is used to merge relations and enhance the semantic understanding of the rule body by focusing on important relations.
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (6)
Here, $Q$, $K$, and $V$ represent queries, keys, and values, respectively, in the attention mechanism, and $d_{k}$ is the dimensionality of the keys. The softmax function is applied to the scaled dot-product of $Q$ and $K$ (scaled by $\sqrt{d_{k}}$), ensuring that the attention weights sum to one. This mechanism allows the model to focus on the most relevant parts of the rule body, facilitating better integration of relational semantics.
The objective function is used to train the model by optimizing the semantic consistency between the rule body and the rule head.
@@
In this equation, $Z$ is a set of rule pairs, $R$ is the set of all relations, $v_{r_{h}}$ is a one-hot vector indicating the rule head, and $\theta_{z}$ represents the learned probabilities associated with each rule. The model maximizes the likelihood of the correct rule head given the rule body, effectively learning high-quality logical rules that capture deep semantic relationships in the knowledge graph.
@@ 4 Methodology
We aim to leverage human bias in sequential decision-making and manipulation problems without compromising interpretability. Human-constructed environments and tasks introduce this bias, represented in our formulation as Responsibility Sharing (RS). Household environments often contain auxiliary objects that, while not mandatory, can be used to indirectly manipulate target objects. RS allows for the division of the original problem into sub-problems through shared responsibilities. For example, an agent can hold an object directly or use a bag to hold the object, with the bag sharing the responsibility of holding the object. By using RS, the agent can simplify complex tasks by breaking them down into more manageable sub-tasks involving auxiliary objects, thereby improving the agent’s effectiveness by leveraging inductive bias in human-constructed problem settings.
To decide when to utilize these auxiliary objects, we train an interpretable model, Optimized Rule Synthesis (ORS), which generates rules for the appropriate conditions under which the agent should delegate responsibilities to these objects. To ensure robustness, we use Counterfactual Plan Generation (CPG) to create a dataset demonstrating how auxiliary objects impact performance changes. This approach ensures that the agent uses auxiliary objects only when it leads to performance improvements. Together, these components enhance the effectiveness and efficiency of TAMP formulation in an interpretable manner. The following subsections elaborate on these components and the problem setup adopted in this study.
@@ 4.1 Problem Setup
We consider fully observable domains with a symbolic state space defined under first-order logic $S$, configuration space $X$, action space $A$, and deterministic transition function $\delta:S\times A\rightarrow S$. We assume a finite set of goals $G$. Each $g\in G$ is a binary condition function $g:S\rightarrow\{0,1\}$ indicating if a symbolic state is a goal state $S_{\text{goal}}(g)$ or $s_{g}$. The goal is to reach a goal state $s_{g}$ starting from an initial state $s_{0}$ following a task plan $P=\{a_{1},a_{2},...,a_{T}\}$, where $T$ is the required number of actions to reach the goal.
Following prior TAMP formulation  [3, 4, 5, 68], we employ a search among the symbolic state space $S$, starting from the initial state $s_{0}$ to find a task plan. This search can either be guided or unguided. IRS as a heuristic is orthogonal to both of these approaches, as it can guide the search by dividing the problem into sub-problems independent of the underlying exploration strategy. In this study, we used a logically unguided search to find task plans. In our formulation, we adopt the concept of responsibility sharing, where the agent can utilize auxiliary and non-mandatory objects if it is inferred to increase the agent’s effectiveness in completing the task. To utilize them, the original problem is divided into sub-problems $\Psi$, which define the usage of the auxiliary objects (e.g., first carry objects on the tray, then carry the tray to the target position). These sub-problems are defined under the set of available actions, $A_{\psi}\subset A$, and the set of sub-goal states, $S_{\text{goal}}(g_{\psi})$. Hence, the original objective remains, but with additional requirements to satisfy each sub-goal state where auxiliary objects are utilized.
To determine when to use auxiliary objects, we propose a rule-based decision-making mechanism, ORS. ORS is trained with a dataset $D$ prepared by the CPG process. This dataset contains initial and goal state descriptions in terms of first-order logic, such as tray on table or picked gripper obj, and a label whether using the auxiliary objects would increase the agent’s performance. ORS is trained with multiple tasks simultaneously and generates a set of rules $S_{c}$, defining appropriate conditions to use auxiliary objects in an environment.
IRS uses these rules to determine when to employ responsibility sharing as a heuristic, thereby improving the agent’s effectiveness and interpretability in decision-making by leveraging human bias when it is inferred to be beneficial.
Figure 3: Counterfactual Plan Generation (CPG) for Dataset Construction
@@ 4.2 Counterfactual Plan Generation
Counterfactual Plan Generation (CPG) aims to generate sequential task plans that utilize the auxiliary objects available in the environment (Figure 3). This method involves three main steps:
@@ 4.2.1 Generating a Feasible Plan
Sequential decision-making and manipulation problems are defined by logical states, actions, goals, and physical constraints. Following the TAMP formulation, the agent employs a search in the state space following a transition function, $s^{\prime}=\delta(s,a)$. Starting from the initial state $s_{0}$, the search algorithm expands from this state by taking actions $a_{t}$, transitioning to the next state $s^{\prime}$, and gradually reaching a goal state $s_{g}$. This generates a series of actions following the formulation detailed in Equation 1, resulting in a task plan, $P=\{a_{1},a_{2},...,a_{T}\}$, where $T$ represents the number of actions required to reach the goal state from the initial state. CPG utilizes Multi-Bound Tree Search  [65], as detailed in Section 3.2, as the search algorithm for finding a task plan due to its domain-agnostic and geometrically aware nature, making it well-suited for these types of problems.
@@ 4.2.2 Generating a Counterfactual Plan
Next, we generate a counterfactual plan using the real plan that was generated in the previous step. The aim of this counterfactual generation is to determine how the plan would change if we utilized the auxiliary objects, i.e., if the agent shared the responsibility. Since the real plan is generated by the uninformed search, it does not utilize these objects as they increase the depth of the solution. Using the same initial state $s_{0}$, transition function $\delta$, and search method (which is uniformed), we introduce an additional sub-goal $s_{i}$, that satisfies the responsibility sharing conditions (e.g., if a tray will be used to carry the objects, they are first placed on the tray). The search expands from the initial state $s_{0}$ until it reaches the goal state $s_{g}$ while satisfying the intermediate state $s_{i}$. This generates a counterfactual series of actions, forming a counterfactual task plan, $P^{\prime}=\{a_{1},a_{2},...,a_{T^{\prime}}\}$, where $T^{\prime}$ represents the number of actions required to reach the goal state from the initial state while employing Responsibility Sharing.
@@ 4.2.3 Dataset Construction
Figure 4: Optimized Rule Synthesis (ORS) for Generating Responsibility Sharing (RS) Conditions
@@ 4.3 Optimized Rule Synthesis (ORS)
Optimized Rule Synthesis (ORS) aims to generate rules for an embodied agent, defining the appropriate conditions for responsibility sharing with auxiliary objects present in the environment (Figure 4). To enhance overall effectiveness and interpretability, ORS integrates Rule-based Representative Learner (RRL)  [25] and Correlation and Order-Aware Rule Learning (CARL)  [26] (Section 3.4 and 3.5). The integration logic involves parsing rules from various sources into a structured format, generating all possible combinations of these rules, and evaluating their applicability to data samples. This process includes calculating the confidence levels for each rule combination and selecting the best-performing rules.
In the integrated approach, the rules are evaluated against the initial and goal states, $s_{0}$ and $s_{g}$, to determine the appropriate conditions for responsibility sharing, taking into account all tasks and environments. The system iteratively checks each combination against the data to find the best match, balancing accuracy and interpretability. The selected set of rules is consistent across different tasks, increasing the generalizability of ORS compared to informed search approaches.
ORS checks whether $s_{0}$ and $s_{g}$ meet all specified conditions of a rule using logical operations like AND or OR. Based on whether the conditions are met, a predicted label (indicating the usage of auxiliary objects, as described in Section 4.2.3) is assigned to each data sample, accompanied by a confidence score representing the proportion of conditions satisfied:
Confidence=Number of conditions satisfiedTotal number of $\text{Confidence}=\frac{\text{Number of conditions satisfied}}{\text{Total number of conditions}}$ (8)
Two key factors in this integration are accuracy and interpretability. Accuracy refers to the correctness of the rule-based decisions, while interpretability relates to how easily these rules can be understood and applied by humans. There is often a trade-off between these two factors: highly accurate models may be complex and difficult to interpret, while simpler, more interpretable models might not achieve the same level of accuracy  [69]. Accuracy can be measured using the formula:
Accuracy=True Positives+True NegativesTotal number of $\text{Accuracy}=\frac{\text{True Positives}+\text{True Negatives}}{\text{Totalnumber of samples}}$ (9)
Interpretability is often inversely related to the complexity of the rules and can be simplified as the number of conditions in the rules.
The integration logic ensures a balance between accuracy and interpretability. Using a validation set, the model optimizes both the accuracy and interpretability of the rules. This balance is achieved through an iterative process of testing and refining the rules, ultimately leading to a set of high-quality, interpretable rules that can effectively handle the usage of auxiliary objects in complex household tasks. ORS generates all possible combinations by combining rule components based on logical operations (AND and OR), starting from a minimum threshold (minimum number of conditions) up to the total available conditions. To optimize the rule-based system, ORS iteratively tests increasing numbers of minimum thresholds for combining conditions, seeking the best Balance Score between prediction accuracy and rule interpretability, as defined in Equation 10.

### v2

@@ 3.2 Multi-Bound Tree Search
The discrete elements of the LGP formulation give rise to a decision tree populated with sequences of symbolic states. The feasibility of each pathway is assessed by solving the NLP produced by the state sequence. To mitigate the high computational cost of solving these NLPs, Multi-Bound Tree Search (MBTS) initially resolves relaxed versions of Equation 1 to prune the search space  [58]. For this work, we utilize two bounds: $P_{seq}$ (joint optimization of mode-switches) and $P_{path}$ (full motion planning).
@@ 3.3 Counterfactual Scenarios & Individual Treatment Effect (ITE)
We quantify the utility of auxiliary objects using the Potential Outcomes framework [43, 66]. Let $W\in\{0,1\}$ denote the binary treatment variable, where $W=1$ indicates the application of Responsibility Sharing (utilizing auxiliary objects) and $W=0$ indicates the baseline strategy (direct manipulation). Let $Y\in\mathbb{R}$ represent the outcome of interest, defined as the agent’s performance cost (e.g., total displacement effort). For a specific task instance $i$, the Individual Treatment Effect (ITE) is defined as the difference between the potential outcomes under the treated and control conditions:
@@
In standard observational studies, one of these outcomes is typically unobserved (counterfactual). However, given the deterministic transition dynamics of our simulation, we can explicitly generate both potential outcomes for a fixed initial state $s_{0}$. We generate the factual plan $P$ (standard search, $W=0$) and the counterfactual plan $P^{\prime}$ (constrained sub-goal search, $W=1$) to directly compute $\text{ITE}_{i}$. This value determines the ground-truth label for training: a negative treatment effect ($\text{ITE}_{i}<0$) implies a reduction in cost, indicating that the use of the auxiliary object yields a positive performance gain.
@@ 3.4 Rule-Based Representation Learner (RRL)
To extract interpretable logical conditions from high-dimensional continuous data, we employ the Rule-Based Representation Learner (RRL) [63]. Traditional decision trees often struggle with scalability on large continuous datasets. RRL addresses this by merging the interpretability of discrete rules with the efficiency of gradient-based optimization. The model utilizes a binarization layer to discretize continuous features and employs Gradient Grafting to bridge the gap between discrete logic and continuous optimization. The update rule is defined as:
@@
This mechanism allows RRL to leverage scalable gradient descent for training while outputting the crisp, discrete logical rules required for symbolic planning.
@@ 3.5 Correlation and Order-Aware Rule Learning (CARL)
While RRL efficiently processes flat feature vectors, TAMP environments are inherently relational (e.g., spatial topologies). To capture these dependencies, we utilize Correlation and Order-Aware Rule Learning (CARL) [29]. Unlike traditional methods that treat rule bodies as unordered sets of atoms, CARL explicitly models the semantic order of relations within a path to predict the most likely rule head.
To distinguish the semantic role of a relation based on its position in the rule chain (e.g., first vs. last condition), the model augments relation embeddings with sinusoidal positional encodings. These encoded inputs are then processed via attention mechanisms to enforce semantic consistency.
Functionally, CARL employs the attention mechanism in two complementary phases to turn symbolic sequences into rich semantic representations.
First, in the Correlation Module, a multi-head self-attention layer is applied over all relations to estimate a global correlation matrix. An active-relations selector uses this to choose only the most informative relations as queries. This prunes the search space, reducing the complexity of relation interactions and focusing resources on relevant candidates.
Second, in the Semantic Learner, multi-head cross-attention is used to recursively merge adjacent relations in the rule body. Here, the query $Q$ represents the current rule-body embedding, while the keys $K$ and values $V$ represent candidate relations. The attention weights effectively determine which relations are semantically compatible with the current rule context, compressing the rule body while enforcing logical consistency.
The model minimizes the cross-entropy loss over the set of training triples $Z$ to maximize the likelihood of the correct rule head $r_{h}$ given the rule body $r_{b}$:
@@
where $v_{r_{h}}$ is the one-hot ground truth vector for the head relation. The term $\theta_{z}\in\mathbb{R}^{|R|+1}$ represents the predicted probability distribution over possible rule heads (including a null head) derived from the attention outputs. Consequently, $\theta_{z}$ serves as the confidence score during rule extraction, ensuring that selected rules are both logically valid and semantically consistent with the underlying knowledge graph.
@@ 4 Methodology
We aim to leverage human-centric environmental bias in sequential decision-making and manipulation problems without compromising interpretability. To achieve this, we formulate Responsibility Sharing (RS) as a mechanism where auxiliary objects share the task load with the agent—for instance, using a bag to hold objects during transport rather than manipulating them directly. The following subsections detail the mathematical formulation of the problem (Section 4.1) and the three core components of our framework: dataset generation (CPG), rule synthesis (ORS), and the heuristic planner (IRS).
@@ 4.1 Problem Setup
We consider fully observable domains with a symbolic state space defined under first-order logic $S$, configuration space $X$, action space $A$, and deterministic transition function $\delta:S\times A\rightarrow S$. We assume a finite set of goals $G$. Each $g\in G$ is a binary condition function $g:S\rightarrow\{0,1\}$ indicating if a symbolic state is a goal state $S_{\text{goal}}(g)$ or $s_{g}$. The goal is to reach a goal state $s_{g}$ starting from an initial state $s_{0}$ following a task plan $P=\{a_{1},a_{2},...,a_{T}\}$, where $T$ is the required number of actions.
Following prior TAMP formulation  [19, 60, 21, 57], we employ a search among the symbolic state space $S$. While this search is typically guided by heuristics, IRS acts as an orthogonal meta-heuristic: it guides the search by structurally dividing the problem into sub-problems, independent of the underlying exploration strategy. In this study, we utilize a logically unguided search to isolate and validate the performance gains provided specifically by the IRS formulation.
In our formulation, we adopt the concept of responsibility sharing, where the agent can utilize auxiliary and non-mandatory objects if inferred to increase effectiveness. To efficiently handle the combinatorial complexity of this choice, we decompose the TAMP problem into a set of specialized sub-problems, $\Psi$. This design choice allows us to isolate the high-level decision (whether to use a tray) from the low-level motion planning, treating Responsibility Sharing as a distinct meta-heuristic. The original problem is divided into sub-problems $\Psi$, which define the usage of the auxiliary objects (e.g., first carry objects on the tray, then carry the tray to the target position). These sub-problems are defined under the set of available actions, $A_{\psi}\subset A$, and the set of sub-goal states, $S_{\text{goal}}(g_{\psi})$. Hence, the original objective remains, but with additional requirements to satisfy each sub-goal state where auxiliary objects are utilized.
@@ 4.2 Counterfactual Plan Generation
To train our rule learner, we require a dataset that objectively quantifies the physical benefit of using auxiliary objects. Since manual human labeling is costly and subjective, we introduce Counterfactual Plan Generation (CPG) to automatically extract ground-truth labels by systematically comparing alternative execution strategies. CPG aims to generate sequential task plans that utilize the auxiliary objects available in the environment (Figure 3). This method involves three main steps:
Figure 3: Counterfactual Plan Generation (CPG) for Dataset Construction. The process compares a standard Real Plan ($P$) against a Counterfactual Plan ($P^{\prime}$) that specifically enforces the use of auxiliary objects. By calculating the Individual Treatment Effect (ITE) based on the cost difference between $P$ and $P^{\prime}$, we determine the ground-truth label for each scenario (Positive if auxiliary usage reduces cost, Negative otherwise), resulting in the labeled dataset $D$ used to train ORS.
@@ 4.2.1 Generating a Feasible Plan
The first step in calculating the treatment effect is to establish a baseline performance metric (the control group). We achieve this by generating a feasible plan using standard uninformed search, which represents the agent’s default behavior when it does not explicitly prioritize responsibility sharing. Sequential decision-making and manipulation problems are defined by logical states, actions, goals, and physical constraints. Following the TAMP formulation, the agent employs a search in the state space following a transition function, $s^{\prime}=\delta(s,a)$. Starting from the initial state $s_{0}$, the search algorithm expands from this state by taking actions $a_{t}$, transitioning to the next state $s^{\prime}$, and gradually reaching a goal state $s_{g}$. This generates a series of actions resulting in a task plan, $P=\{a_{1},a_{2},...,a_{T}\}$. CPG utilizes Multi-Bound Tree Search  [58] as the search algorithm for finding a task plan due to its domain-agnostic and geometrically aware nature.
@@ 4.2.2 Generating a Counterfactual Plan
Next, we generate a counterfactual plan using the real plan that was generated in the previous step. The aim of this counterfactual generation is to determine how the plan would change if we utilized the auxiliary objects. Since the real plan is generated by the uninformed search, it does not utilize these objects as they increase the depth of the solution. Using the same initial state $s_{0}$, transition function $\delta$, and search method (which is uninformed), we introduce an additional sub-goal $s_{i}$, that satisfies the responsibility sharing conditions (e.g., if a tray will be used to carry the objects, they are first placed on the tray). The search expands from the initial state $s_{0}$ until it reaches the goal state $s_{g}$ while satisfying the intermediate state $s_{i}$. This generates a counterfactual series of actions, forming a counterfactual task plan, $P^{\prime}=\{a_{1},a_{2},...,a_{T^{\prime}}\}$.
@@ 4.2.3 Dataset Construction
To ensure the validity of this labeling process and rule out potential skewness caused by the uninformed search (which prioritizes plan depth), we conducted a validation study on the serving task (detailed in Section 5.1). We compared our dataset labels against a planner-independent geometric heuristic that calculates the optimal $L^{2}$ path cost for both direct (individual transport) and auxiliary-based (shared responsibility) strategies. The results showed a strong alignment between our planner-derived labels and the geometric ground truth. This confirms that our dataset accurately reflects the intrinsic importance of object usage based on environmental structure, rather than exploiting the disadvantages of the uninformed search algorithm.
Figure 4: The Optimized Rule Synthesis (ORS). The framework takes the labeled dataset $D$ as input and employs two complementary generators: RRL (extracting structural logic) and CARL (extracting relational dependencies). The resulting candidate rules are pooled and iteratively selected via a bounded greedy search that optimizes a Balance Score, weighing accuracy against interpretability. The final output is a stable, concise set of conditions ($S_{c}$) that dictates when Responsibility Sharing should be applied.
@@ 4.3 Optimized Rule Synthesis (ORS)
Optimized Rule Synthesis (ORS) is designed to learn interpretable rules that specify precisely when the agent should employ Responsibility Sharing (RS) with auxiliary objects. The method operates on a supervised dataset $D=\{(x_{i},y_{i})\}_{i=1}^{N}$, constructed via Counterfactual Plan Generation (CPG). Here, $x_{i}=\Phi(s_{0}^{i},g^{i})\in\mathbb{R}^{d}$ represents the feature encoding of the $i$-th initial–goal state pair, and $y_{i}\in\{-1,+1\}$ indicates the ground-truth label: +1+1 if using RS is beneficial, and 1-1 otherwise.
The feature encoding function $\Phi(s,g)$ maps the symbolic state $s$ and logical goal $g$ into a fixed-size binary feature vector $x\in\{0,1\}^{d}$. This vector consists of flattened ground predicates representing the environment’s configuration and task constraints. Specifically, the feature space $\Phi$ encompasses:
  • 1.
    Spatial Configuration: Boolean indicators for the position of all manipulable and static objects relative to discrete locations (e.g., $obj_{i}\_on\_table_{north}$, $tray\_on\_table_{south}$).
  • 2.
    Goal Constraints: Boolean flags specifying the target conditions for the task (e.g., $goal\_at\_table_{east}$).
  • 3.
    Action & Context Primitives: Indicators for environmental capabilities (e.g., is_helper_existis\_helper\_exist) and the set of high-level abstract actions valid for the current task domain (e.g., $action_{pour}$, $action_{handover}$, $action_{pick}$).
For RRL, this binary vector is used directly as input. For CARL, these active predicates are projected into a knowledge graph structure, where entities (e.g., objects, tables) are nodes and their spatial or functional relationships serve as edges.
To ensure robustness and coverage, ORS acts as an ensemble selector. It integrates two distinct rule-learning architectures as complementary generators: Rule-based Representation Learner (RRL) [63], which constructs logical rules from discretized continuous features, and Correlation and Order-Aware Rule Learning (CARL) [29], which extracts rules from a knowledge graph, preserving correlation and order information. Rather than using these models directly, ORS aggregates their outputs and selects a sparse, high-performing subset of rules that balances accuracy with interpretability.
4.3.1 Joint Rule Pool Generation
To achieve a robust rule set, ORS employs a dual-generator strategy that leverages the distinct strengths of two architectures. We utilize RRL to capture strict structural constraints (e.g., explicit negations and specific object placements) from the feature vectors, while CARL is employed to extract complex relational dependencies and semantic chains from the knowledge graph. In the initial phase, both RRL and CARL are employed to generate a comprehensive set of candidate rules from the dataset $D$. Let $\mathcal{R}^{\text{CARL}}$ and $\mathcal{R}^{\text{RRL}}$ denote the sets of rules extracted by CARL and RRL, respectively. These are combined into a joint rule pool $\mathcal{R}=\mathcal{R}^{\text{CARL}}\cup\mathcal{R}^{\text{RRL}}$, which is subsequently deduplicated to remove semantically equivalent rules sharing the same antecedent and label.
Each candidate rule $r\in\mathcal{R}$ is defined by two components: a Boolean condition $\varphi_{r}:\mathbb{R}^{d}\to\{0,1\}$ (e.g., tray is present $\land$ object count $\geq 3$) evaluated on input $x_{i}$, and a label $\ell(r)\in\{-1,+1\}$ indicating the recommended decision when the condition holds.
4.3.2 Rule-Set Classification
ORS functions as a meta-learner that selects a subset $S\subseteq\mathcal{R}$ to form an interpretable classifier. This classifier maps feature vectors $x$ to RS decisions $\hat{y}$ via a voting mechanism. For a single rule $r$, the prediction on instance $x_{i}$ is determined by its condition satisfaction:
$\hat{y}_{i}^{(r)}=\begin{cases}\ell(r),&\text{if }\varphi_{r}(x_{i})=1,\\ -\ell(r),&\text{if }\varphi_{r}(x_{i})=0.\end{cases}$ (5)
This formulation implies that a rule predicts its assigned label if satisfied, and the opposite label otherwise. The agreement of a rule with the ground truth is given by the indicator $\gamma_{i}(r)=\mathbb{I}[\hat{y}_{i}^{(r)}=y_{i}]$, and the individual accuracy of a rule $r$ over the dataset is defined as $\mathrm{Acc}(r;D)=\frac{1}{N}\sum_{i=1}^{N}\gamma_{i}(r)$.
To aggregate predictions from the selected subset $S$, we employ a signed voting scheme. The ensemble score is calculated as $s_{S}(x)=\sum_{r\in S}\varphi_{r}(x)\,\ell(r)$. The final classification decision $h_{S}(x)$ is then derived by thresholding this sum:
$h_{S}(x)=\begin{cases}+1,&\text{if }s_{S}(x)>0,\\ -1,&\text{otherwise}.\end{cases}$ (6)
The performance of the rule set is evaluated on a validation split $D_{\mathrm{val}}$ using the standard accuracy metric $\mathrm{Acc}(S;D_{\mathrm{val}})=\frac{1}{|D_{\mathrm{val}}|}\sum_{i}\mathbb{I}[h_{S}(x_{i})=y_{i}]$.
4.3.3 Interpretability and Balanced Score
In the domain of Task and Motion Planning (TAMP), we define interpretability not as sparsity (minimizing conditions), but as informational sufficiency or descriptiveness. While general machine learning often prefers shorter rules to prevent overfitting, robotic decision-making requires rules that capture sufficient physical context to justify an action to a human observer and grounded task execution [56, 33].
For instance, a sparse rule such as Positive $\longleftarrow$ $obj_{south}$ recommends an action based solely on object location, failing to explain the physical mechanism enabling the task. In contrast, a longer, context-aware rule such as Positive $\longleftarrow$ $obj_{south}$ $\land$ helper exists $\land$ $\neg goal_{south}$ explicitly identifies the necessary preconditions (e.g., the existence of a tray/helper) that make the action valid. To quantify this preference for context, we define the interpretability of a rule set $S$, denoted $\mathcal{I}(S)$, as the average rule length normalized by the maximum possible conditions $L_{\max}$:
$\mathcal{I}(S)=\frac{1}{|S|}\sum_{r\in S}\frac{\text{length}(r)}{L_{\max}}$ (7)
where $\text{length}(r)$ is the number of atomic conditions in the rule antecedent.
However, maximizing length arbitrarily can obscure the causal explanation through the dilution effect [38]. To distinguish between necessary causal context and arbitrary or delusional verbosity, we utilize a Balance Score that jointly optimizes Accuracy and Interpretability:
$\mathrm{Score}(S;D_{\mathrm{val}})=\alpha\cdot\mathrm{Acc}(S;D_{\mathrm{val}})+(1-\alpha)\cdot\mathcal{I}(S)$ (8)
By maximizing this combined objective, the system favors rules that are sufficiently descriptive (high $\mathcal{I}(S)$) to be intelligible. Crucially, the accuracy term ($\mathrm{Acc}$) acts as a semantic filter: conditions that increase length without contributing to (or maintaining) the predictive validity of the rule are rejected, ensuring that the retained predicates represent meaningful physical constraints rather than noise.
4.3.4 Iterative Search Procedure
Finding the optimal subset of rules that maximizes the Balanced Score is a combinatorial problem. Since enumerating all subsets is intractable for a realistic rule pool size $|\mathcal{R}|$, ORS employs a bounded greedy search strategy. This process is controlled by two hyperparameters: $\beta$, the number of random repetitions, and $K_{\max}$, the maximum number of rules allowed in set $S$.
To ensure generalization and reduce variance, the search procedure is repeated $\beta$ times (with $\beta=10$ in our experiments). In each iteration, the dataset $D$ is randomly partitioned into training ($D_{\mathrm{train}}$), validation ($D_{\mathrm{val}}$), and test ($D_{\mathrm{test}}$) subsets. The candidate generators, CARL and RRL, are trained exclusively on $D_{\mathrm{train}}$ to populate the rule pool $\mathcal{R}^{(\mathrm{iter})}$.
The selection process begins with an empty set $S=\emptyset$. In each step, the algorithm greedily adds the rule $r\in\mathcal{R}^{(\mathrm{iter})}\setminus S$ that yields the largest increase in $\mathrm{Score}(S;D_{\mathrm{val}})$. This strictly relies on the validation set to guide selection, preventing data leakage from the test set. The process terminates when no remaining rule improves the score or when $|S|$ reaches $K_{\max}$. After $\beta$ repetitions, the rules selected most frequently across all runs are aggregated to form the final, stable condition set $S_{c}$. The complete procedure is detailed in Algorithm 1.
Algorithm 1 Optimized Rule Synthesis (ORS)
1:Input: Dataset $D=\{(x_{i},y_{i})\}_{i=1}^{N}$
2:Hyperparameters: $\beta$ (repetitions), $K_{\max}$ (max rules), $\alpha$
3:Output: Final RS rule set $S_{c}$